# Synthesizing Sentiment-Controlled Feedback For Multimodal Text and Image Data

### v3

@@ 3.1.2. Preprocessing
The CMFeed dataset consists of multiple images per sample, corresponding news text, post likes and shares, and human comments along with reactions and shares. The comments for each post have been sorted based on Facebook’s ‘most-relevant’ criterion, which prioritizes the comments with the highest likes and shares. The comments have been preprocessed as follows: emoticons converted to words using the Demoji library (Solomon, [n. d.]); blank comments removed; contractions expanded; special and accented characters eliminated; punctuations and numbers purged to reduce noise; stop-words removed; and comments converted to lowercase. These steps are designed to reduce noise while retaining key sentiment cues in the text. This approach is based on the understanding that core sentiments are primarily conveyed through words, not punctuation or numbers. As per Facebook’s ethical protocols, we used manual data scraping and strategically chose $1000$ posts from each of four news handles—BBC News, Sky News, Fox News, and NY Daily News—to promote diversity and reduce bias. Initially, $4000$ posts were collected; however, after preprocessing, $3646$ posts remained. Subsequently, all associated comments were processed, yielding a total of $61734$ comments. As summarized in Table 2, on average, each news post received $65.1$ likes, while comments averaged $10.5$ likes. The average length of news text was $655$ words. Additionally, each post contained an average of $3.7$ images.
Table 2. CMFeed dataset’s overview.
@@
3.1.3. Annotation Strategy
: To determine the ground-truth sentiment labels for the comments, we obtained sentiment scores using four pre-trained models: FLAIR Library (Akbik et al., 2019), SentimentR (Rinker, 2017), DistilBERT (Sanh et al., 2019), and RoBERTa (Liu et al., 2019). Each of these models has unique capabilities: FLAIR specializes in capturing contextual variations in text using a neural network approach. SentimentR is designed to analyze textual sentiments by evaluating linguistic cues within the text. DistilBERT and RoBERTa are both transformer-based models optimized for understanding the nuances of language through self-attention mechanisms. DistilBERT offers a lighter, faster variant of BERT that retains most of its predictive power, whereas RoBERTa is trained on an even larger corpus with more robust fine-tuning, enhancing its ability to discern complex sentiment patterns.
We adopted a majority voting strategy for annotation, similar to that used in the construction of the IEMOCAP dataset (Busso et al., 2008), retaining data samples that received the same sentiment class from at least three of the four models. The remaining samples were excluded and marked with the sentiment class ‘XX.’ The models generated sentiment values of $-1$ or $1$ along with a confidence score. We calculated a score by multiplying the sentiment value by its confidence, then normalized this score to a range between 00 (negative sentiment) and $1$ (positive sentiment). Averages of these normalized scores across the models provided the final sentiment labels. To further ensure robustness, we also employed a safety margin from $0.49$ and $0.51$ and marked the labels with the score in this range as ‘XX.’ Of the total $61734$ samples analyzed, $57222$ met the criteria for inclusion in further experiments based on the above filtering strategies, ensuring high confidence in the ground-truth labels.
The sentiment prediction was conducted on comments to capture and analyze direct human emotional reactions, independent of the original text or imagery. This approach aims to understand user reactions, which are direct, spontaneous, and personal, thus providing insights into user sentiments. The objective is to enable the feedback synthesis system to emulate this human-like directness and spontaneity. To ensure the accuracy of the constructed ground-truth sentiment labels, a human evaluation was conducted. $50$ evaluators ($25$ males and $25$ females, average age $30\pm 2.73$ years) assessed the sentiment of $50$ randomly selected pairs of input image and text. The results showed that $90.88$% of evaluators (standard deviation $7.59$%) agreed on the consistency between the expressed sentiment and the assigned sentiment label.
Table 3. Representative samples from the CMFeed dataset. Here, ‘PLikes’ (Post Likes) and ‘CLikes’ (Comment Likes) show the number of likes for the post and comment, respectively; ‘Shares’ denotes number of shares for the post and ‘C’ represents comment’s sentiment class ($1$: positive, 00: negative).
@@ 3.1.3. Annotation Strategy
If it leaks then it’s easy to clean. Bin strikes: The people using baths and hiring skips to store rubbish, Helen Sikora has been keeping rubbish bags in her bath so she can easily clean up any leaks. Edinburgh residents have told how they have hired skips and used bathtubs to store rubbish, as waste piles up … [Uncaptioned image] 464 29 Dump it on streets if they do not bother to collect it. 29 0
@@
The CMFeed dataset’s sample instances have been depicted in Table 3. Multiple images and comments correspond to each news post, enabling the feedback synthesis model to learn the comments’ contextual diversity and relevance with the input. The CMFeed dataset can be accessed at zenodo.org/records/11409612 along with detailed description of its metadata, collection process, intended usage, licensing, and handling. The corresponding code can be accessed at github.com/MIntelligence-Group/CMFeed/.
3.2. Task Formulation
@@
Given an environment $E=\{T,I_{1},I_{2},\ldots,I_{n}\}$, where $T$ denotes input text and $I_{1},I_{2},\ldots,I_{n}$ denote $n$ input images. Each image is comprised of $m$ objects $o_{1},o_{2},\ldots,o_{m}$, while the text $T$ is made up of a dictionary of $k$ words $w_{1},w_{2},\ldots,w_{k}$. The task is to generate a feedback towards environment $E$ under a sentiment-controlled manner where ‘sentiment-controlled’ implies the feedback aligns with a specified sentiment $S$, with $S$ being either 00 for a negative sentiment or $1$ for a positive sentiment.
3.3. Proposed Feedback Synthesis System
The proposed system, as illustrated in Fig. 1, includes two networks for processing textual and visual data, each with an encoder, a decoder, and a control layer. It extracts textual and visual features using a text transformer (Vaswani et al., 2017) and a Faster R-CNN model (Ren et al., 2015), respectively. These features are then used by the decoders to generate feedback. The system also includes a similarity module to evaluate the feedback’s contextual alignment with human comments. From input images and text, the proposed system uses encoders to extract features, integrating visual and textual information to comprehend the content and its context thoroughly. It also learns from comments to analyze sentiment and user engagement, which helps in crafting responses that mirror human emotional reactions. Overall, the system synthesizes these insights to generate sentiment-controlled feedback that is contextually relevant and emotionally resonant, adapting its output to match the desired sentiment.
3.3.1. Textual Encoder
@@
The textual encoder utilizes a transformer model (Vaswani et al., 2017), featuring global encoding and textual attention mechanisms. Global encoding is enhanced by a convolution-gated unit to improve textual representation, reduce repetition, and maintain semantic relevance. The textual attention component includes multi-headed self-attention, comprising a self-attention layer and a feed-forward layer. Positional embedding captures token positioning, and normalization finalizes the process to produce the textual context vector $z_{t}^{*}$. The feed-forward network (FFN) includes input and output layers with a dimension of $512$ and a hidden layer of $2048$. The FFN’s output for a specific input $x$ is defined in Eq.1.
(1) $FFN(x)=max(0,xW_{1}+b_{1})W_{2}+b_{2}$
@@
3.3.2. Visual Encoder
The top three images from each sample are inputted into a visual encoder, using blank images when fewer are available. Features extracted from these images are concatenated to form a visual context vector $z_{i}^{*}$. This process employs a pre-trained Faster R-CNN model (Ren et al., 2015). According to Eq. 2, CNN layers produce feature maps that the Region Proposal Network (RPN) uses to create anchor boxes with binary scores based on Intersection Over Union (IoU) values (Rosebrock, 2016). These anchors are then classified and regressed to yield classified boxes. A total of 1601 classes are assigned to these boxes, and their features are combined into a global feature vector. Faster R-CNN is selected for its efficiency and precision in detecting small and varied objects, suitable for non-real-time settings. The choice to use the top three images strikes a balance between retaining essential visual content and minimizing blank images in posts with fewer images, considering most posts have at least three images.
(2) $Objectiveness\ Score=\begin{cases}Positive;\hskip 14.45377ptIoU>0.7\\ Negative;\hskip 10.84006ptIoU<0.3\\ No\ score;\hskip 10.84006pt0.3<IoU<0.7\\ \end{cases}$
Figure 1. Proposed feedback synthesis system’s architecture with encoder, decoder, and controllability blocks for textual and visual data. The decoder, convolution-gated unit, control layer, and similarity modules are shown as distinct subblocks while off & on neurons are shown as black & white circles.
3.3.3. Attention
@@
(3) $\begin{split}&\ \ \ z=Attention(Q,K,V)=softmax(\frac{Q.K^{T}}{\sqrt{d_{k}}})V\ \end{split}$
where $Q$, $K$, and $V$ represent matrices that contain all queries, keys, and values, respectively, with $d_{k}$ as the scaling factor and $K^{T}$ denoting the transpose of $K$. To achieve a comprehensive subspace representation, the mechanism computes multiple attention heads using distinct sets of Query, Key, and Value matrices. The queries, keys, and values undergo projection $head$ times, resulting in multiple attention heads $h_{1},h_{2},...,h_{head}$, where $head$ signifies the total number of heads. These heads are then concatenated and multiplied by the weight matrix $W$, producing the intermediate output vector $z^{\prime}$, as delineated in Eq. 4.
(4) $\begin{split}&h_{i}=Attention(QW^{Q}{i},KW^{K}{i},VW^{V}{i})\\ &z^{\prime}=Concat(h_{1},h_{2}...,h_{head})W^{O}\ \end{split}$
@@ 3.3.4. Decoder
The textual and visual decoders share a similar structure, each consisting of two main blocks: a self-attention block and an encoder-decoder attention block, enhanced with positional encoding and normalization for improved efficiency and accuracy. The textual decoder processes the textual context vector $z_{t}$, while the visual decoder handles the visual context vector $z_{i}$. Both decoders also receive the ground-truth comment as input. The self-attention layer uses future position masking to focus only on previous positions in the output sequence, with its weight matrices initially random and refined during training. The encoder-decoder attention layer uses the context vectors $z_{t}$ or $z_{i}$ as keys and values, respectively. Late fusion is applied via concatenation to maintain distinct image feature information, and a gated convolution unit is incorporated for textual features to reduce repetition in the feedback.
3.3.5. Control Layer
@@
The control layer, positioned after the decoder and before feedback generation, introduces perturbations to ensure feedback aligns with desired sentiments. It utilizes two masks, one each for positive and negative sentiments, altering the input vector via element-wise multiplication as per Eq. 5. This layer functions like a modified dropout layer, selectively activating or deactivating neurons to tune sentiment in the feedback, ensuring it matches the targeted tone.
(5) $O=\begin{cases}mask_{1}*I;\hskip 25.29494ptSentiment=0\\ mask_{2}*I;\hskip 25.29494ptSentiment=1\\ \end{cases}$
@@
where $O$ and $I$ denote the output and input vectors, respectively. In this setup, each mask blocks $x\%$ of neurons, targeting different neuron sets. Consequently, $(100-2x)\%$ of neurons are trained on both sentiments, while $x\%$ are specialized for a specific sentiment, with $x$set at $10$%. This configuration helps direct the feedback towards the desired sentiment tone. During inference, to generate sentiment-specific feedback, neurons trained for the contrasting sentiment are deactivated. For instance, to produce positive feedback, neurons associated with negative sentiment are turned off, and vice versa. This method is crucial for controlling the output sentiment independently of the input text’s sentiment, focusing on steering the generated sentence’s sentiment.
3.3.6. Similarity Module
@@
The similarity module quantitatively assesses the semantic similarity between the feedbacks generated by the proposed system and human comments, using a pre-trained Sentence-BERT (SBERT) model (Reimers and Gurevych, 2019). It transforms individual comments and feedbacks into vectors in an $n$-dimensional embedding space, where $n$ is the size of the embeddings produced by the model. Each dimension represents a distinct linguistic feature or attribute, effectively capturing the sentence’s semantic characteristics. After generating these embeddings, the cosine similarity between the vector representations of the generated feedback and the human-provided comment is computed. This cosine similarity score, a robust metric for this purpose, captures the orientation of the sentence vectors, reflecting their semantic similarity and providing a measure of relevance between the feedbacks and comments.
3.3.7. Interpretability
@@
This section proposes an interpretability technique using the K-Average Additive exPlanation (KAAP) method (Kumar et al., 2022) to analyze the contribution of textual and visual features towards feedback generation. It is based on Shapley Additive exPlanations (SHAP) that is an approximation of Shapley values (Shapley, 1953). As depicted in Fig. 2, it has been incorporated to assess the influence of each visual and textual feature on the sentiment portrayed by the generated feedback. We hypothesize that varying sentiments produced from identical inputs (text + images) should reflect in differential feature importance. It is expected that key features will differ for negative versus positive sentiments. When identical inputs are processed to portray varied sentiments, the model should adjust its focus across different image and text segments, thereby validating our controllability hypothesis.
@@
Figure 2. Representation of the proposed interpretability technique. Here $k_{i}$ and $k_{t}$ denote the number of partitions for image and text, $w_{i}$ is the image’s width, and $L_{t}$ is the text feature vector’s length.
@@
(6) $\mathscr{S}_{i}(f)=\sum_{S\subseteq F\setminus\{i\}}\frac{|S|!(|F|-|S|-1)!}{|F|!}[f(S\cup\{i\})-f(S)]$
@@
(7) $X=[X_{1},X_{2},...,X_{k}],\text{where }X_{i}\subseteq X\text{and}\bigcup_{i=1}^{k}X_{i}=X$
@@
(8) $\mathscr{S}_{\{f_{1}\}}+\mathscr{S}_{\{f_{2}\}}=\mathscr{S}_{\{f_{1},f_{2}\}}-\mathscr{S}_{\{\text{null}\}}$
@@
(9) $\text{KAAP}_{i}=\frac{1}{k}\sum_{j=1}^{k}\mathscr{S}_{i}(X_{j})$
@@
4. Experimental Results
4.1. Training Strategy and Parameter Tuning
The proposed models have been trained for $60$ epochs on an Nvidia V100100100100 GPU, employing $5$-fold cross-validation and an $80$%-$20$% training-testing split.
  • General parameters – Batch size: $16$, learning rate: $0.001$, network optimizer: Adam, loss function: cross-entropy loss, activation function: ReLU.
  • Parameters for the transformer model – Encoder embedding dimensions: $100$, decoder embedding dimensions: $100$, encoder hidden units dimensions: $128$, decoder hidden units dimensions: $128$, encoder dropout: $0.1$, decoder dropout: $0.1$, encoder no. of layers and attention heads: $3$ and $8$, decoder no. of layers & attention heads: $3$ and $8$, metric: accuracy.
  • Parameters for the Faster R-CNN model – No. of epochs: $18$, no. of proposals: $18$, no. of anchor-box classes: $1601$, network optimizer: adaDelta, metric: mAP (mean Average Precision).
4.2. Evaluation Metrics

### v4

@@ III-A2 Preprocessing
The CMFeed dataset includes multiple images per sample, corresponding news text, post likes and shares, and human comments with reactions and shares. Comments are sorted by Facebook’s ‘most-relevant’ criterion, prioritizing the highest likes and shares. Preprocessing converts emoticons to words with Demoji [58], removes blank comments, expands contractions, eliminates special and accented characters, purges punctuation and numbers, removes stop-words, and lowercases text. It reduces noise while retaining key sentiment cues (core sentiments are conveyed mainly through words rather than punctuation or numbers). As per Facebook’s ethical protocols, we used manual scraping and selected $1000$ posts from each of four news handles. Initially $4000$ posts were collected; after preprocessing, $3646$ remained, with all associated comments processed, yielding $61734$ comments. As summarized in Table II, each news post received on average $65.1$ likes, comments averaged $10.5$ likes, news text averaged $655$ words, and each post contained on average $3.7$ images.
TABLE II: CMFeed dataset’s overview.
@@
TABLE III: Representative samples from the CMFeed dataset. Here, ‘PLikes’ and ‘CLikes’ are likes on the post and the comment, ‘Shares’ is the post’s share count and ‘C’ is the comment’s sentiment class (1 positive, 0 negative).
@@
III-A3 Annotation Strategy
To determine the ground-truth sentiment labels for the comments, we obtained sentiment scores using four pre-trained models: FLAIR Library [59], SentimentR [60], DistilBERT [61], and RoBERTa [62]. Each of these models has unique capabilities: FLAIR specializes in capturing contextual variations in text using a neural network approach. SentimentR is designed to analyze textual sentiments by evaluating linguistic cues within the text. DistilBERT and RoBERTa are both transformer-based models optimized for understanding the nuances of language through self-attention mechanisms. DistilBERT offers a lighter, faster variant of BERT that retains most of its predictive power, whereas RoBERTa is trained on an even larger corpus with more robust fine-tuning, enhancing its ability to discern complex sentiment patterns.
We adopted a majority voting strategy for annotation, as in IEMOCAP [63], retaining samples that received the same sentiment class from at least three of four models; the rest were excluded and marked ‘XX.’ The models output sentiment values of 1-1 or $1$ with a confidence score. We computed a score by multiplying sentiment and confidence, normalized it to $[0,1]$ with 0 negative and $1$ positive, and averaged the normalized scores across models to obtain final labels. To ensure robustness, we used a safety margin from $0.49$ to $0.51$ and marked labels in this range as ‘XX.’ $57222$ of $61734$ samples met the criteria, providing high-confidence ground-truth labels.
The sentiment prediction was conducted on comments to capture and analyze direct human emotional reactions, independent of the original text or imagery. This approach aims to understand user reactions, which are direct, spontaneous, and personal, thus providing insights into user sentiments. The objective is to enable the feedback synthesis system to emulate this human-like directness and spontaneity. To ensure the accuracy of the constructed ground-truth sentiment labels, a human evaluation was conducted. $50$ evaluators ($25$ males and $25$ females, average age $30\pm 2.73$ years) assessed the sentiment of $50$ randomly selected pairs of input image and text. The results showed that 90.8890.88% of evaluators (standard deviation 7.597.59%) agreed on the consistency between the expressed sentiment and the assigned sentiment label.
@@ III-A3 Annotation Strategy
The CMFeed dataset’s samples have been depicted in Table III. Multiple images and comments correspond to each news post, enabling the feedback synthesis model to learn the comments’ contextual diversity and relevance to the input. The CMFeed dataset and the corresponding code can be accessed at zenodo.org/records/11409612 and github.com/MIntelligence-Group/CMFeed/, respectively. More details about the CMFeed dataset are provided in Supplementary Material Sec. S2.
III-B Task Formulation
@@
Given an environment $E=\{T,I_{1},I_{2},\ldots,I_{n}\}$, where $T$ denotes input text and $I_{1},I_{2},\ldots,I_{n}$ denote $n$ input images. Each image comprises $m$ objects $o_{1},o_{2},\ldots,o_{m}$, while the text $T$ is made up of a dictionary of $k$ words $w_{1},w_{2},\ldots,w_{k}$. The task is to generate a feedback towards environment $E$ in a sentiment-controlled manner where ‘sentiment-controlled’ implies the feedback aligns with a specified sentiment $S$, with $S$ being either 0 or $1$ for a negative or positive sentiment.
III-C Proposed Feedback Synthesis System
The proposed system (Fig. 1) includes two networks for textual and visual data, each with an encoder, a decoder, and a control layer. The system generates context-aware feedback matching the desired sentiment whose alignment with human comments is checked by the similarity module.
Figure 1: Proposed system’s architecture with encoder, decoder and controllability blocks for textual and visual data. Decoder, convolution-gated unit, control layer, and similarity modules appear as subblocks; on/off neurons as black/white circles.
III-C1 Textual Encoder
@@
The textual encoder utilizes a transformer model [6], featuring global encoding and textual attention mechanisms. Global encoding is enhanced by a convolution-gated unit to improve textual representation, reduce repetition, and maintain semantic relevance. The textual attention component includes multi-headed self-attention, comprising a self-attention layer and a feed-forward layer. Positional embedding captures token positioning, and normalization finalizes the process to produce the textual context vector $z_{t}^{*}$. The feed-forward network (FFN) includes input and output layers with a dimension of $512$ and a hidden layer of $2048$. The FFN’s output for a specific input $x$ is defined in Eq.1.
$FFN(x)=max(0,xW_{1}+b_{1})W_{2}+b_{2}\$(1)
@@
III-C2 Visual Encoder
The top three images from each sample are input to a visual encoder, using blank images when fewer are available; their features are concatenated to form the visual context vector $z_{i}^{*}$. We employ a pre-trained Faster R-CNN [7]: CNN layers yield feature maps that the RPN converts to anchor boxes with binary scores based on IoU [64], which are then classified and regressed into $1601$ classes, whose features are combined into a global vector. Faster R-CNN is selected for efficiency and precision in detecting small and varied objects in non-real-time settings. Using the top three images balances retaining essential visual content and minimizing blanks, since most posts have at least three images.
[…]
III-C3 Attention
@@
$\begin{split}&\ \ \ z=Attention(Q,K,V)=softmax(\frac{Q.K^{T}}{\sqrt{d_{k}}})V\ \end{split}$ (3)
where $Q$, $K$, and $V$ are matrices containing all queries, keys, and values, with $d_{k}$ as scaling factor and $K^{T}$ transpose of $K$. To achieve a comprehensive subspace representation, the mechanism computes multiple attention heads using distinct Query, Key, and Value projections. The queries, keys, and values undergo projection headhead times, yielding heads $h_{1},h_{2},...,h_{head}$, where headhead is the total number of heads. The heads are then concatenated and multiplied by the weight matrix $W$, producing the intermediate output vector $z^{\prime}$, as in Eq. 4.
$\begin{split}&h_{i}=Attention(QW^{Q}{i},KW^{K}{i},VW^{V}{i})\\ &z^{\prime}=Concat(h_{1},h_{2}...,h_{head})W^{O}\ \end{split}$ (4)
@@ III-C4 Decoder
The textual and visual decoders share a similar structure: two blocks (self-attention and encoder–decoder attention) with positional encoding and normalization. The textual decoder processes $z_{t}$, the visual decoder handles $z_{i}$; both also receive the ground-truth comment. The self-attention layer uses future position masking to attend only to previous outputs. The encoder–decoder attention uses $z_{t}$ or $z_{i}$ as keys and values. Late fusion via concatenation preserves distinct image features, and a gated convolution unit is added for textual features to reduce repetition.
III-C5 Control Layer
@@
The control layer, positioned after the decoder and before feedback generation, introduces perturbations to ensure feedback aligns with desired sentiments. It utilizes two masks, one each for positive and negative sentiments, altering the input vector via element-wise multiplication as per Eq. 5. This layer functions like a modified dropout layer, selectively activating or deactivating neurons to tune sentiment in the feedback, ensuring it matches the targeted tone. From a human–machine systems perspective, providing interpretable control signals improves user understanding and trust in the loop. Our mask-based control layer follows this principle by including an inspectable intervention on the decoder state [65].
[…]
@@
where $O$ and $I$ denote the output and input vectors, respectively. Each mask blocks $x\%$ neurons, targeting different neuron sets; consequently, $(100-2x)\%$ of neurons are trained on both sentiments, while $x\%$ are specialized, with $x=10\%$. This directs the feedback toward the desired sentiment. At inference, to generate sentiment-specific feedback, neurons trained for the contrasting sentiment are deactivated; for instance, to produce positive feedback, neurons associated with negative sentiment are turned off, and vice versa. The output sentiment is controlled independently of the input text’s sentiment, steering the generated sentence’s sentiment.
III-C6 Similarity Module
@@
The similarity module quantitatively assesses semantic similarity between generated feedbacks and human comments using a pre-trained Sentence-BERT (SBERT) model [66]. It transforms comments and feedbacks into vectors in an $n$-dimensional embedding space (where $n$ is the model’s embedding size), with each dimension representing a distinct linguistic feature capturing sentence characteristics. Cosine similarity is then computed between the feedback and comment vectors. This score captures the orientation of the sentence vectors, reflecting semantic similarity and providing a relevance measure between feedbacks and comments.
III-C7 Interpretability
@@
This section proposes an interpretability technique using the K-Average Additive exPlanation (KAAP) method [67] to analyze the contribution of textual and visual features toward feedback generation. It is based on an approximation of Shapley values [68]. As depicted in Fig. 2, it assesses the influence of each visual and textual feature on the sentiment portrayed by the generated feedback. We hypothesize that varying sentiments produced from identical inputs (text + images) should reflect in differential feature importance, with key features differing for negative versus positive sentiments. When identical inputs portray varied sentiments, the model should adjust its focus across image and text segments, thereby validating our controllability hypothesis.
@@
Figure 2: Depiction of the proposed interpretability technique. Here $k_{i}$ and $k_{t}$ are number of partitions for image and text, $w_{i}$ is the image’s width and $L_{t}$ is the text feature vector’s length.
@@
$\mathscr{S}_{i}(f)=\sum_{S\subseteq F\setminus\{i\}}\frac{|S|!(|F|-|S|-1)!}{|F|!}[f(S\cup\{i\})-f(S)]$ (6)
@@
$X=[X_{1},X_{2},...,X_{k}],\text{where }X_{i}\subseteq X\text{and}\bigcup_{i=1}^{k}X_{i}=X$ (7)
@@
$\mathscr{S}_{\{f_{1}\}}+\mathscr{S}_{\{f_{2}\}}=\mathscr{S}_{\{f_{1},f_{2}\}}-\mathscr{S}_{\{\text{null}\}}$ (8)
@@
$\text{KAAP}_{i}=\frac{1}{k}\sum_{j=1}^{k}\mathscr{S}_{i}(X_{j})$ (9)
@@
IV Experimental Results
IV-A Training Strategy and Parameter Tuning
The proposed models were trained for $60$ epochs on an Nvidia V100100 GPU using $5$-fold cross-validation with an 80802020 train–test split. General parameters included a batch size $16$, learning rate $0.001$, Adam optimizer, cross-entropy loss, and ReLU activation. For the transformer, encoder and decoder embeddings were $100$, hidden units $128$ (encoder and decoder), dropout $0.1$ (encoder and decoder), with $3$ layers and $8$ attention heads for both encoder and decoder; the metric was accuracy. For Faster R-CNN, training used $18$ epochs, $18$ proposals, $1601$ anchor-box classes, the AdaDelta optimizer, and mean Average Precision as the metric.
IV-B Evaluation Metrics